# Laude: LLM-Assisted Unit Test Generation and Debugging of Hardware DEsigns

### v1

@@ 5.3 Debuggability Analysis of Unit Tests
For Gemini-2.5 Pro, the DA distributions closely mirror the distributions observed in Section˜5.2. Both NLS and NLSC achieve high DA values within 80% – 100% for BC08 – BC10. However, for BC01 – BC07, DA drops to intermediate bins despite non-zero AR, revealing cases where bugs are detected but insufficiently propagated through the design to produce sustained divergence. This shows that successful attacks do not guarantee strongly divergent failure traces. For NL, DeepSeek R1 remains ineffective while NLS improves performance but suffers from high variance, suggesting unstable reasoning when only buggy code is provided. In contrast, NLSC consistently produces high DA scores for complex bug targets, demonstrating that few-shot guidance enables the model to detect bugs and propagate their effects through multiple simulation cycles. This experiment shows that generating a unit test for real-world debugging requires creativity that can be challenging for state-of-the-art LLMs.
Figure 9: Divergent Attack (DA) Distribution for DeepSeek R1 using the same format as Figure˜8. NL is predominantly concentrated in Bin 1 (0%–20%) across buggy codes. The NLS shows mixed results with high variance. NLSC demonstrates the strongest performance, particularly for BC08 – BC10.
Figure 10: (10(a)): Debugging success rates of NL, NLS, and NLSC for various LLMs for buggy combinational circuit design codes.  (10(b)): Debugging success rates of NL, NLS, and NLSC for various LLMs for buggy sequential circuit design codes.
@@ 5.4 Effectiveness of Iterative Debugging
For NL, Gemini-2.5 Pro exhibit high success rates (median $\approx 0.85$ for combinational and $\approx 0.73$ for sequential). In contrast, DeepSeek R1 remains highly unstable; its distribution is notably wide and dispersed (median $\approx 0.5$), indicating a lack of consistent reasoning patterns when the model is forced to debug with unit tests generated via NL. In NLS, Gemini-2.5 Pro and DeepSeek R1 result in higher performance. The availability of the source code acts as a critical grounding mechanism for the open-source models. DeepSeek R1 achieve the largest relative gains, with medians reaching between 0.7 and 0.8. This suggests that while these models lack the intuition to generate high-quality unit tests in NL, their reasoning engines are fully capable of analyzing specific bugs when the design circuit is exposed. The NLSC provides a universal, though more incremental refinement for both LLMs. The NLSC pushes the median success rates for all models above 0.75 for combinatorial, closing the gap between the proprietary and the open-source models. Gemini-2.5 Pro retains a consistent reasoning as evidenced by tighter distributions near the upper bound of the metric compared to the larger variance in the DeepSeek R1.
@@ 7 Conclusion
Laude provides a scalable and automated solution for localizing, diagnosing, and debugging functional errors in the hardware design source code using unit tests. The experimental results on a commonly used benchmark and a variety of functional bugs demonstrate its effectivity in detecting and debugging functional errors. Laude demonstrated that while the current close and open-source LLMs can reason on complex hardware verification tasks, there are broad scopes for improvement.
@@ 8 Limitations
Figure 11: Attack Rate (AR) Distribution for Gemini-2.5 Pro. This figure applies the established 10×10 glyph heatmap format with 5-bin sparklines to analyze the performance of Gemini 2.5 Pro. The $10\times 10$ gymph heatmap with sparklines utilize equally spaced 5-bin scale: Bin 1 (0%–20%) through Bin 5 (80%–100%) from left to right. Results: Unlike DeepSeek R1, Gemini-2.5 Pro demonstrates high robustness in the zero-shot NL setting (dominant green, median $80\%$). However, the NLS configuration introduces unexpected instability, particularly for targets BC02 and BC07 where median AR drops to the 40%–60% range, with a specific collapse ($20\%$) for Source BC09 on Target BC07. The few-shot NLSC configuration effectively restores consistent high sensitivity across all pairs.
Figure˜11 shows the AR Sparkline plots for Gemini-2.5 Pro using the NL, NLS, and NLSC configurations, while Figure˜12 shows the AR results for DeepSeek R1. These plots corresponds to the AR Sparkline Plot analysis of Section˜5.1.
Figure 12: Attack Rate (AR) Distribution for DeepSeek R1 using the same format as Figure˜11. The model largely fails in the zero-shot NL setting (dominant red, median $20\%$) but achieves near-universal robustness (dominant green, median $80\%$) once structural context is introduced in NLS and NLSC.
@@
Figure 13: Attack Rate (AR) Distribution for Gemini Flash. Gemini-2.5 Flash maintains high sensitivity ($90\%$, dominant green) in the zero-shot NL. The NLS and NLSC configurations further improve the results, pushing the median AR to $\approx 100\%$ across all targets
Figure 14: Attack Rate (AR) Distribution for Qwen-2.5 Coder. The NL configuration shows a comprehensive failure (dominant red, median $20\%$) across nearly all source-target pairs. However, the introduction of the system prompt in NLS delivers better AR results, restoring high sensitivity (dominant green, $80\%$) for BC08–BC10.
@@ B.1 Bug Sensitivity Analysis of Unit Tests
Figure 15: Divergence Rate (DR) Distribution for Gemini-2.5 Flash. We observe that Gemini-2.5 Flash exhibits a concerning rigidity in the NLS setting; its distributions remain largely static compared to the NL for the complex targets (BC01—-BC07).
Figure 16: Divergence Rate (DR) Distribution for Qwen-2.5 Coder. The sparklines utilize the 5-bin scale: Bin 1 (0%–20%) through Bin 5 (80%–100%). Results: NL is predominantly concentrated in Bin 1 (0%–20%) across buggy codes. The NLS and NLSC demonstrates the strongest performance, particularly for BC08 – BC10.
Figures˜13 and 14 show the AR sparkline results. For Qwen, NL is largely ineffective, with AR distributions consistently stagnating in the lowest quartile ($20\%$). However, the transition to NLS effectively restores performance, shifting the distributions into the $80\%--100\%$ range for BC08–BC10, a trend that stabilizes further with NLSC. In contrast, Gemini-2.5 Flash exhibits exceptional robustness from the outset. Unlike Qwen, it does not suffer from zero-shot collapse; both NL and NLS maintain high sensitivity with median ARs consistently $90\%$. The NLSC configuration serves only to saturate this already high performance, pushing the AR distributions to $\approx 100\%$. This experiment highlights a distinct architectural divergence: while Qwen shows a more gradual capability curve, Gemini-2.5 Flash demonstrates robustness, achieving higher results in all configurations.
@@ B.2 Bug Specificity Analysis of Unit Tests
For NL, Qwen-2.5 Coder performs the worst characterized by distributions spreading towards the lowest bin (0% - 20%). Gemini-2.5 Flash performs better, achieving the highest results for BC08 – BC10. Moving to the NLS configuration, Qwen demonstrates a robust architectural recovery, shifting its DR results toward the upper quartiles ($50\%$) for almost all bins, proving its capability to utilize instructions to correct its generation logic. In stark contrast, Gemini-2.5 Flash exhibits a concerning rigidity in the NLS setting; its distributions remain largely static compared to the NL for the complex targets (BC01–-BC07). It is only under the NLSC configuration that both models converge to saturation, but the trajectory differs: Qwen arrives there via steady instruction-following improvements, while Gemini relies entirely on the few-shot context to unlock its high-efficacy states. This experiment shows that while Qwen-2.5 Coder progressively improves its reasoning via instruction following, Gemini-2.5 Flash’s detection capability remains latent until unlocked by few-shot examples, exposing Gemini’s critical dependency on in-context learning grounding.
@@ B.3 Debuggability Analysis of Unit Tests
For Gemini-2.5 Flash, the DA distributions closely mirror the distributions observed in Section˜B.2. Specifically, the NLS configuration fails to yield a substantial improvement over NL, with the majority of targets (particularly BC01 – BC07) remaining trapped in the lower efficacy bins ($40\%$). This indicates that simply providing the buggy code is insufficient for the Flash model to achieve satisfactory results, as done by the Pro model. In contrast, Qwen-2.5 Coder demonstrates a decisive shift in the NLS configuration, where the DA distributions migrate significantly toward the higher bins ($60\%-100\%$). Both models eventually converge to the highest DA results in the NLSC configuration. The disparity in the NLS configuration confirms the results found in the previous analysis in Sections˜B.2 and B.1
@@
Figure 19: (19(a)): Debugging success rates of NL, NLS, and NLSC for various LLMs for buggy combinational circuit design codes.  (19(b)): Debugging success rates of NL, NLS, and NLSC for Gemini-2.5 Flash and Qwen-2.5 32B Coder for buggy sequential circuit design codes.
@@ B.4 Effectiveness of Iterative Debugging
The results reveal a striking performance disparity between the models, contradicting the trends observed in testbench generation. For combinatorial circuits, Gemini-2.5 Flash shows high robustness when debugging, maintaining a consistently high mean success rate across all configurations: $\mu=0.84$ in NL, $\mu=0.87$ in NLS, and $\mu=0.89$ in NLSC, which is the most effective configuration. This suggests that for this iterative repair task, Gemini Flash possesses a stable intrinsic debugging capability that is largely independent of the prompt’s context level. On the other hand, Qwen-2.5 Coder shows significant sensitivity and, compared to Gemini Flash, lower performance. In the NL configuration, it struggles to debug effectively ($\mu=0.13$), likely due to the lack of context. We can see that providing the faulty code (NLS) yields a substantial improvement to $\mu=0.72$. Within NLSC configuration, it achieves the best performances overall, but these results are still, on average, 14% lower than what Gemini Flash can achieve. The relative performances of the two LLMs don’t significantly change when focusing on sequential circuits. Due to their inherently higher complexity of sequential, the debugging success rates are lower compared to combinatorial, but Gemini-2.5 Flash still manages to achieve an average success rate $80\%$ for any configuration. Qwen performs very poorly on NL, as the distribution is centered towards lower percentages, but manages to gain an average of 44% when shifting to NLS configuration.
@@ C.2 Effectiveness of Iterative Debugging using Violin Plots
Figure 20: (20(a)): Debugging success rates of NL, NLS, and NLSC for Gemini-2.5 Pro and DeepSeek R1 for buggy combinational circuit design codes.  (20(b)): Debugging success rates of NL, NLS, and NLSC for Gemini 2.5 Pro and DeepSeek R1 for buggy sequential circuit design codes.
Figure˜20(a) and Figure˜20(b) show the violin plot distributions for the success rates for Gemini-2.5 Pro and DeepSeek R1 across the three configurations. For combinational designs, we can see that for Gemini-2.5 Pro, the violin shapes are top-heavy across all configurations. These upper sections indicate a ceiling effect – the model consistently averages a high score and hits the maximum possible success rate with very few outliers. In contrast, DeepSeek R1 in the NL configuration shows a flat, stretched shape, meaning its performance is effectively random—sometimes good, often bad. Adding prompts (NLS and NLSC) compresses this shape into a tighter blob in the upper range, showing that structure improves the score and stabilizes the model’s behavior.
For sequential designs, we see that Gemini-2.5 Pro remains consistent; the violins are compact and narrow, showing little variation between different runs. The interesting detail appears in DeepSeek R1. In the NL setting, the mass is concentrated at the bottom (consistent failure). While the NLS prompt moves this mass clearly upward, the NLSC (few-shot) configuration results in a wider, fatter shape. This visible expansion in width indicates that while few-shot examples help some test cases, they introduce higher variance in others, making the model less predictable than with simple instructions.
@@
Figure 21: (21(a)): Debugging success rates of NL, NLS, and NLSC for Gemini-2.5 Flash and Qwen-2.5 32B Coder for buggy combinational circuit design codes.  (21(b)): Debugging success rates of NL, NLS, and NLSC for Gemini-2.5 Flash and Qwen-2.5 32B Coder for buggy sequential circuit design codes.
@@
We extend our analysis to Gemini-2.5 Flash and Qwen-2.5 Coder using violin plots to visualize their stability and responsiveness to different prompting strategies. Considering the combinational circuits, Figure˜21(a) reveals a contrast in baseline capability. Gemini-2.5 Flash is remarkably consistent; even in the simple NL configuration, the distribution is concentrated high up ($\mu=0.84$), and adding context in NLSC only slightly tightens this top-heavy shape ($\mu=0.89$). It works well “out of the box.” In contrast, Qwen-2.5 Coder is effectively broken in the NL setting, with a flat distribution stuck at the bottom ($\mu=0.13$). However, the violin shape changes dramatically with the NLS prompt, jumping massively to a median of $\mu=0.72$. This “gap” proves that Qwen possesses the logic to solve these circuits but is completely dependent on specific instructions to unlock it, whereas Gemini-2.5 Flash is self-sufficient.
For sequential designs, Figure˜21(b) highlights a critical weakness in Qwen’s handling of context. Gemini-2.5 Flash remains “prompt-invariant”; its violins are almost identical across all three settings ($\mu\approx 0.81$), showing that neither instructions nor few-shot examples significantly change its stable behavior. Qwen-2.5 Coder, however, shows a unique regression. While it improves from NL ($\mu=0.25$) to NLS ($\mu=0.69$), adding the few-shot examples in NLSC actually hurts performance, dropping the median to $\mu=0.57$. The violin shape for NLSC becomes wider and shifts lower than NLS, indicating that for this model, the extra text from the few-shot examples acts as a distraction rather than a guide.
The violin plots observed in this experiment expose a “Contextual Distraction” effect in Qwen-2.5 Coder. While Gemini-2.5 Flash ignores the prompt format and relies on its internal reasoning (staying flat around $\mu=0.81$), Qwen is highly sensitive. Crucially, in sequential tasks, “more is less” – simple instructions (NLS) work best, while the longer context of few-shot learning (NLSC) confuses the model, causing a performance drop of 12%. This suggests smaller open-source models may struggle to attend to relevant details when the prompt becomes too long. This comparison highlights a trade-off. Gemini-2.5 Flash is robust, delivering consistent high scores regardless of the prompt method. In contrast, Qwen-2.5 Coder is more sensitive: it fails without instructions (NL) and becomes better with NLS and NLSC.

### v2

@@ 5.3 Debuggability Analysis of Unit Tests
For Gemini-2.5 Pro, the DA distributions closely mirror the distributions observed in Section˜5.2. Both NLSC@0 and NLSC@5 achieve high DA values within 80% – 100% for BC08 – BC10. However, for BC01 – BC07, DA drops to intermediate bins despite non-zero AR, revealing cases where bugs are detected but insufficiently propagated through the design to produce sustained divergence. This shows that successful attacks do not guarantee strongly divergent failure traces. For NL, DeepSeek R1 remains ineffective while NLSC@0 improves performance but suffers from high variance, suggesting unstable reasoning when only buggy code is provided. In contrast, NLSC@5 consistently produces high DA scores for complex bug targets, demonstrating that few-shot guidance enables the model to detect bugs and propagate their effects through multiple simulation cycles. This experiment shows that generating a unit test for real-world debugging requires creativity that can be challenging for state-of-the-art LLMs.
Figure 9: Divergent Attack (DA) Distribution for DeepSeek R1 using the same format as Figure˜8. NLS@0 is predominantly concentrated in Bin 1 (0%–20%) across buggy codes. The NLSC@0 shows mixed results with high variance. NLSC@5 demonstrates the strongest performance, particularly for BC08 – BC10.
Figure 10: (10(a)): Debugging success rates of NLS@0, NLSC@0, and NLSC@5 for various LLMs for buggy combinational circuit design codes.  (10(b)): Debugging success rates of NLS@0, NLSC@0, and NLSC@5 for various LLMs for buggy sequential circuit design codes.
@@ 5.4 Effectiveness of Iterative Debugging
For NLS@0, Gemini-2.5 Pro exhibit high success rates (median $\approx 0.85$ for combinational and $\approx 0.73$ for sequential). In contrast, DeepSeek R1 remains highly unstable; its distribution is notably wide and dispersed (median $\approx 0.5$), indicating a lack of consistent reasoning patterns when the model is forced to debug with unit tests generated via NLS@0. In NLSC@0, Gemini-2.5 Pro and DeepSeek R1 result in higher performance. The availability of the source code acts as a critical grounding mechanism for the open-source models. DeepSeek R1 achieve the largest relative gains, with medians reaching between 0.7 and 0.8. This suggests that while these models lack the intuition to generate high-quality unit tests in NLS@0, their reasoning engines are fully capable of analyzing specific bugs when the design circuit is exposed. The NLSC@5 provides a universal, though more incremental refinement for both LLMs. The NLSC@5 pushes the median success rates for all models above 0.75 for combinatorial, closing the gap between the proprietary and the open-source models. Gemini-2.5 Pro retains a consistent reasoning as evidenced by tighter distributions near the upper bound of the metric compared to the larger variance in the DeepSeek R1.
@@ 7 Conclusion
Laude provides a scalable and automated solution for localizing, diagnosing, and debugging functional errors in the hardware design source code using unit tests. The experimental results on a commonly used benchmark and a variety of functional bugs demonstrate its effectivity in detecting and debugging functional errors. Laude demonstrated that while the current closed- and open-source LLMs can reason on complex hardware verification tasks, there are broad scopes for improvement.
@@ 8 Limitations
Figure 11: Attack Rate (AR) Distribution for Gemini-2.5 Pro. This figure applies the established 10×10 glyph heatmap format with 5-bin sparklines to analyze the performance of Gemini 2.5 Pro. The $10\times 10$ gymph heatmap with sparklines utilize equally spaced 5-bin scale: Bin 1 (0%–20%) through Bin 5 (80%–100%) from left to right. Results: Unlike DeepSeek R1, Gemini-2.5 Pro demonstrates high robustness in the zero-shot NLS@0 setting (dominant green, median $80\%$). However, the NLSC@5 configuration introduces unexpected instability, particularly for targets BC02 and BC07 where median AR drops to the 40%–60% range, with a specific collapse ($20\%$) for Source BC09 on Target BC07. The few-shot NLSC@5 configuration effectively restores consistent high sensitivity across all pairs.
Figure˜11 shows the AR Sparkline plots for Gemini-2.5 Pro using the NLS@0, NLSC@0, and NLSC@5 configurations, while Figure˜12 shows the AR results for DeepSeek R1. These plots corresponds to the AR Sparkline Plot analysis of Section˜5.1.
Figure 12: Attack Rate (AR) Distribution for DeepSeek R1 using the same format as Figure˜11. The model largely fails in the zero-shot NL setting (dominant red, median $20\%$) but achieves near-universal robustness (dominant green, median $80\%$) once structural context is introduced in NLSC@0 and NLSC@5.
@@
Figure 13: Attack Rate (AR) Distribution for Gemini Flash. Gemini-2.5 Flash maintains high sensitivity ($90\%$, dominant green) in the NLS@0. The NLSC@0 and NLSC@5 configurations further improve the results, pushing the median AR to $\approx 100\%$ across all targets
Figure 14: Attack Rate (AR) Distribution for Qwen-2.5 Coder. The NLS@0 configuration shows a comprehensive failure (dominant red, median $20\%$) across nearly all source-target pairs. However, the introduction of the system prompt in NLSC@0 delivers better AR results, restoring high sensitivity (dominant green, $80\%$) for BC08–BC10.
@@ B.1 Bug Sensitivity Analysis of Unit Tests
Figure 15: Divergence Rate (DR) Distribution for Gemini-2.5 Flash. We observe that Gemini-2.5 Flash exhibits a concerning rigidity in the NLSC@0 setting; its distributions remain largely static compared to the NLS@0 for the complex targets (BC01—-BC07).
Figure 16: Divergence Rate (DR) Distribution for Qwen-2.5 Coder. The sparklines utilize the 5-bin scale: Bin 1 (0%–20%) through Bin 5 (80%–100%). Results: NLS@0 is predominantly concentrated in Bin 1 (0%–20%) across buggy codes. The NLSC@0 and NLSC@5 demonstrates the strongest performance, particularly for BC08 – BC10.
Figures˜13 and 14 show the AR sparkline results. For Qwen, NLS@0 is largely ineffective, with AR distributions consistently stagnating in the lowest quartile ($20\%$). However, the transition to NLSC@0 effectively restores performance, shifting the distributions into the $80\%--100\%$ range for BC08–BC10, a trend that stabilizes further with NLSC@5. In contrast, Gemini-2.5 Flash exhibits exceptional robustness from the outset. Unlike Qwen, it does not suffer from zero-shot collapse; both NLS@0 and NLSC@0 maintain high sensitivity with median ARs consistently $90\%$. The NLSC@5 configuration serves only to saturate this already high performance, pushing the AR distributions to $\approx 100\%$. This experiment highlights a distinct architectural divergence: while Qwen shows a more gradual capability curve, Gemini-2.5 Flash demonstrates robustness, achieving higher results in all configurations.
@@ B.2 Bug Specificity Analysis of Unit Tests
For NLS@0, Qwen-2.5 Coder performs the worst characterized by distributions spreading towards the lowest bin (0% - 20%). Gemini-2.5 Flash performs better, achieving the highest results for BC08 – BC10. Moving to the NLSC@0 configuration, Qwen demonstrates a robust architectural recovery, shifting its DR results toward the upper quartiles ($50\%$) for almost all bins, proving its capability to utilize instructions to correct its generation logic. In stark contrast, Gemini-2.5 Flash exhibits a concerning rigidity in the NLSC@0 setting; its distributions remain largely static compared to the NLS@0 for the complex targets (BC01–-BC07). It is only under the NLSC@5 configuration that both models converge to saturation, but the trajectory differs: Qwen arrives there via steady instruction-following improvements, while Gemini relies entirely on the few-shot context to unlock its high-efficacy states. This experiment shows that while Qwen-2.5 Coder progressively improves its reasoning via instruction following, Gemini-2.5 Flash’s detection capability remains latent until unlocked by few-shot examples, exposing Gemini’s critical dependency on in-context learning grounding.
@@ B.3 Debuggability Analysis of Unit Tests
For Gemini-2.5 Flash, the DA distributions closely mirror the distributions observed in Section˜B.2. Specifically, the NLSC@0 configuration fails to yield a substantial improvement over NLS@0, with the majority of targets (particularly BC01 – BC07) remaining trapped in the lower efficacy bins ($40\%$). This indicates that simply providing the buggy code is insufficient for the Flash model to achieve satisfactory results, as done by the Pro model. In contrast, Qwen-2.5 Coder demonstrates a decisive shift in the NLSC@0 configuration, where the DA distributions migrate significantly toward the higher bins ($60\%-100\%$). Both models eventually converge to the highest DA results in the NLSC@5 configuration. The disparity in the NLSC@0 configuration confirms the results found in the previous analysis in Sections˜B.2 and B.1
@@
Figure 19: (19(a)): Debugging success rates of NLS@0, NLSC@0, and NLSC@5 for various LLMs for buggy combinational circuit design codes.  (19(b)): Debugging success rates of NLS@0, NLSC@0, and NLSC@5 for Gemini-2.5 Flash and Qwen-2.5 32B Coder for buggy sequential circuit design codes.
@@ B.4 Effectiveness of Iterative Debugging
The results reveal a striking performance disparity between the models, contradicting the trends observed in testbench generation. For combinatorial circuits, Gemini-2.5 Flash shows high robustness when debugging, maintaining a consistently high mean success rate across all configurations: $\mu=0.84$ in NLS@0, $\mu=0.87$ in NLSC@0, and $\mu=0.89$ in NLSC@5, which is the most effective configuration. This suggests that for this iterative repair task, Gemini Flash possesses a stable intrinsic debugging capability that is largely independent of the prompt’s context level. On the other hand, Qwen-2.5 Coder shows significant sensitivity and, compared to Gemini Flash, lower performance. In the NLS@0 configuration, it struggles to debug effectively ($\mu=0.13$), likely due to the lack of context. We can see that providing the faulty code (NLSC@0) yields a substantial improvement to $\mu=0.72$. Within NLSC@5 configuration, it achieves the best performances overall, but these results are still, on average, 14% lower than what Gemini Flash can achieve. The relative performances of the two LLMs don’t significantly change when focusing on sequential circuits. Due to their inherently higher complexity of sequential, the debugging success rates are lower compared to combinatorial, but Gemini-2.5 Flash still manages to achieve an average success rate $80\%$ for any configuration. Qwen performs very poorly on NLS@0, as the distribution is centered towards lower percentages, but manages to gain an average of 44% when shifting to NLSC@0 configuration.
@@ C.2 Effectiveness of Iterative Debugging using Violin Plots
Figure 20: (20(a)): Debugging success rates of NLS@0, NLSC@0, and NLSC@5 for Gemini-2.5 Pro and DeepSeek R1 for buggy combinational circuit design codes.  (20(b)): Debugging success rates of NLS@0, NLSC@0, and NLSC@5 for Gemini 2.5 Pro and DeepSeek R1 for buggy sequential circuit design codes.
Figure˜20(a) and Figure˜20(b) show the violin plot distributions for the success rates for Gemini-2.5 Pro and DeepSeek R1 across the three configurations. For combinational designs, we can see that for Gemini-2.5 Pro, the violin shapes are top-heavy across all configurations. These upper sections indicate a ceiling effect – the model consistently averages a high score and hits the maximum possible success rate with very few outliers. In contrast, DeepSeek R1 in the NLS@0 configuration shows a flat, stretched shape, meaning its performance is effectively random—sometimes good, often bad. Adding prompts (NLSC@0 and NLSC@5) compresses this shape into a tighter blob in the upper range, showing that structure improves the score and stabilizes the model’s behavior.
For sequential designs, we see that Gemini-2.5 Pro remains consistent; the violins are compact and narrow, showing little variation between different runs. The interesting detail appears in DeepSeek R1. In the NLS@0 setting, the mass is concentrated at the bottom (consistent failure). While the NLSC@0 prompt moves this mass clearly upward, the NLSC@5 (few-shot) configuration results in a wider, fatter shape. This visible expansion in width indicates that while few-shot examples help some test cases, they introduce higher variance in others, making the model less predictable than with simple instructions.
@@
Figure 21: (21(a)): Debugging success rates of NLS@0, NLSC@0, and NLSC@5 for Gemini-2.5 Flash and Qwen-2.5 32B Coder for buggy combinational circuit design codes.  (21(b)): Debugging success rates of NLS@0, NLSC@0, and NLSC@5 for Gemini-2.5 Flash and Qwen-2.5 32B Coder for buggy sequential circuit design codes.
@@
We extend our analysis to Gemini-2.5 Flash and Qwen-2.5 Coder using violin plots to visualize their stability and responsiveness to different prompting strategies. Considering the combinational circuits, Figure˜21(a) reveals a contrast in baseline capability. Gemini-2.5 Flash is remarkably consistent; even in the simple NLS@0 configuration, the distribution is concentrated high up ($\mu=0.84$), and adding context in NLSC@5 only slightly tightens this top-heavy shape ($\mu=0.89$). It works well “out of the box.” In contrast, Qwen-2.5 Coder is effectively broken in the NLS@0 setting, with a flat distribution stuck at the bottom ($\mu=0.13$). However, the violin shape changes dramatically with the NLSC@0 prompt, jumping massively to a median of $\mu=0.72$. This “gap” proves that Qwen possesses the logic to solve these circuits but is completely dependent on specific instructions to unlock it, whereas Gemini-2.5 Flash is self-sufficient.
For sequential designs, Figure˜21(b) highlights a critical weakness in Qwen’s handling of context. Gemini-2.5 Flash remains “prompt-invariant”; its violins are almost identical across all three settings ($\mu\approx 0.81$), showing that neither instructions nor few-shot examples significantly change its stable behavior. Qwen-2.5 Coder, however, shows a unique regression. While it improves from NLS@0 ($\mu=0.25$) to NLSC@0 ($\mu=0.69$), adding the few-shot examples in NLSC@5 actually hurts performance, dropping the median to $\mu=0.57$. The violin shape for NLSC becomes wider and shifts lower than NLS, indicating that for this model, the extra text from the few-shot examples acts as a distraction rather than a guide.
The violin plots observed in this experiment expose a “Contextual Distraction” effect in Qwen-2.5 Coder. While Gemini-2.5 Flash ignores the prompt format and relies on its internal reasoning (staying flat around $\mu=0.81$), Qwen is highly sensitive. Crucially, in sequential tasks, “more is less” – simple instructions (NLSC@0) work best, while the longer context of few-shot learning (NLSC@5) confuses the model, causing a performance drop of 12%. This suggests smaller open-source models may struggle to attend to relevant details when the prompt becomes too long. This comparison highlights a trade-off. Gemini-2.5 Flash is robust, delivering consistent high scores regardless of the prompt method. In contrast, Qwen-2.5 Coder is more sensitive: it fails without instructions (NLS@0) and becomes better with NLSC@0 and NLSC@5.